OpenStudy (theloshua):

If 30.0 mL of 0.150 M aqueous sodium hydroxide is mixed with 30.0 mL of 0.150 M aqueous hydrochloric acid in a calorimeter at an initial temperature of 25.0 degrees Celsius, what is the enthalpy change of this reaction if the final temperature reached in the calorimeter is 27.5 degrees Celsius? NaOH + HCl yields NaCl + H2O

OpenStudy (aaronq):

For calorimetry reactions, use the equation: \(q=m_{water}*C_p*\Delta T\) where m is the mass of water (or solution in this case) Cp is the specific heat capacity for the system (i would assume it to be the same as water here) \(\Delta T\) is change in temperature the system underwent

OpenStudy (aaronq):

oh yea and q is the heat evolved (which is equal to the change in enthalpy under constant pressure, which I assume it's valid here)

OpenStudy (theloshua):

okayy..... so..... wait

OpenStudy (aaronq):

lol so \(C_p=4.18 ~J/~ ^oC*g\) So we have \(\sf q=m*(4.18 ~J/~ ^oC*g)*(T_{final}-T_{initial})\) we need the mass and the temperatures..

OpenStudy (theloshua):

ummm.... but water isnt involved in the reaction...

OpenStudy (aaronq):

I know, but we're measuring the temperature of the water (solution), and that is how (indirectly) we're getting the change in enthalpy

OpenStudy (theloshua):

but the solution isnt water XD so you dont know its heat capacity....

OpenStudy (theloshua):

so isnt it irrelevant?

OpenStudy (aaronq):

It's common in these problems to make the simplification that the specific heat capacity is equal to that of water, unless you're told otherwise... same with the density of the solution, it's equal to that of water

OpenStudy (theloshua):

okayy.... I hope your right... :p

OpenStudy (aaronq):

Unless you're told (in the question) otherwise, i'm right lol

OpenStudy (theloshua):

okay

OpenStudy (theloshua):

so ummmm hold on let me plug in the equation

OpenStudy (theloshua):

what is the mass of the equation? :/

OpenStudy (aaronq):

the mass of the solution, the density is assumed to be the same as water's so \(\sf mass=density*volume=1~g/mL*60~mL=60~g\)

OpenStudy (theloshua):

im confused right now....

OpenStudy (aaronq):

What are you confused about?

OpenStudy (theloshua):

you took the density and volume of water to find the mass right?

OpenStudy (aaronq):

yep, the definition for density is: \(\sf density=\dfrac{mass}{volume}\) so i just rearranged

OpenStudy (theloshua):

but doesnt the density and volume of water depend on the amount? or no?

OpenStudy (aaronq):

the density depends on a lot of things, most importantly temperature. Again, these problems make the simplification that the density of the solution is the same as that of water .. which is further simplified and rounded to 1 g/mL. The volume is given in the question

OpenStudy (theloshua):

oh, you added the two volumes?

OpenStudy (aaronq):

yeah

OpenStudy (aaronq):

Thats the volume of the solution

OpenStudy (theloshua):

okayyyy so \[Q = 60 \times 4.18 \times (27.5 - 25)\]

OpenStudy (theloshua):

right?

OpenStudy (aaronq):

yep. theres one final thing, because this is the energy absorbed by the water it is an endothermic process, but the reaction is exothermic as energy was released. \(\sf q_{absorbed}=-q_{released}\) you need to change the sign

OpenStudy (theloshua):

oh okay, but what about the values given (0.250 M)?

OpenStudy (theloshua):

where do those come in?

OpenStudy (aaronq):

they don't, unless they want the molar enthalpy of the reaction. You'll come across questions that include more info than you need and you have to know what to use and what is irrelevant

OpenStudy (theloshua):

okay, so hold on

OpenStudy (theloshua):

The answer would be -627?

OpenStudy (aaronq):

yep, just tack on some units to that

OpenStudy (theloshua):

okay, thnx so much!!!!

OpenStudy (aaronq):

no problem !